### v1

# M2SE: A Multistage Multitask Instruction Tuning Strategy for Unified Sentiment and Emotion Analysis

Abstract
Sentiment analysis and emotion recognition are crucial for applications such as human-computer interaction and depression detection. Traditional unimodal methods often fail to capture the complexity of emotional expressions due to conflicting signals from different modalities. Current Multimodal Large Language Models (MLLMs) also face challenges in detecting subtle facial expressions and addressing a wide range of emotion-related tasks. To tackle these issues, we propose M2SE, a Multistage Multitask Sentiment and Emotion Instruction Tuning Strategy for general-purpose MLLMs. It employs a combined approach to train models on tasks such as multimodal sentiment analysis, emotion recognition, facial expression recognition, emotion reason inference, and emotion cause-pair extraction. We also introduce the Emotion Multitask dataset (EMT), a custom dataset that supports these five tasks. Our model, Emotion Universe (EmoVerse), is built on a basic MLLM framework without modifications, yet it achieves substantial improvements across these tasks when trained with the M2SE strategy. Extensive experiments demonstrate that EmoVerse outperforms existing methods, achieving state-of-the-art results in sentiment and emotion tasks. These results highlight the effectiveness of M2SE in enhancing multimodal emotion perception. The dataset and code are available at https://github.com/xiaoyaoxinyi/M2SE.
Index Terms:
Multimodal Large Language models, Multitask, Sentiment, Emotion.
publicationid: pubid: $0$–$00$ 2024 IEEE
I Introduction
Sentiment analysis and emotion recognition are critical components in a variety of domains, including human-computer interaction[1], depression detection[2], and more. These tasks involve identifying sentiments and emotional states in human communication. While unimodal approaches—those that rely on individual modalities such as face[3], text[4], or audio[5]—have achieved notable success in sentiment and emotion detection, they are inherently limited in their ability to fully capture the complexity of human emotions. Each modality alone provides only a partial view of the emotional landscape, making it difficult to analyze data in a comprehensive and holistic manner. Furthermore, unimodal systems are prone to misinterpretations, especially in cases where emotions are expressed indirectly or ambiguously, such as with sarcasm[6], irony, or subtle emotional cues. These challenges highlight the need for more integrated and multimodal approaches to sentiment and emotion analysis that can overcome the limitations of single-modality methods.
Existing methods for fusing multimodal information typically rely on cross-modal interactions to learn from different modalities and ultimately perform sentiment analysis or emotion recognition tasks[7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. However, there is an inherent relationship between sentiment and emotion that these methods often struggle to capture. Simply completing sentiment or emotion tasks in isolation does not effectively model the complex interplay between specific emotions and corresponding sentiments.
Moreover, while current Multimodal Large Language Models (MLLMs) demonstrate strong performance in tasks such as visual language understanding [30, 31, 32], they often underperform in tasks like multimodal sentiment analysis and emotion recognition in zero-shot settings. As a result, fine-tuning remains essential. As shown in Figure 1, although existing MLLMs achieve promising results when fine-tuned for a single task, their performance on other tasks frequently deteriorates. This underscores the need for a general-purpose MLLM capable of performing well across a variety of sentiment- and emotion-related tasks, and capturing both sentiment and emotion in a more integrated and comprehensive manner.
Figure 1: A comparison of MLLMs trained solely on the MSA task, solely on the ER task, and using the M2SE strategy. (a) In the MSA task, the model trained on MSA performs well, while the model trained on ER performs poorly. (b) In the ER task, the model trained on ER performs well, while the model trained on MSA performs poorly. (c) EmoVerse, trained with the M2SE strategy, not only excels on both MSA and ER tasks but also successfully performs the emotion reason inference task.
To fine-tune any MLLM with multitask capabilities, we explore the relationships between different tasks and propose a novel strategy called M2SE: Multistage Multitask Sentiment and Emotion Instruction Tuning Strategy for common MLLMs. Through our exploration of task relationships, M2SE applies a multistage approach, where different tasks are introduced at each stage of training, and each task is assigned a different sampling rate based on its importance. By using this strategy, we train Emotion Universe (EmoVerse), a model based on traditional MLLM architecture that integrates both visual and textual modalities. Notably, EmoVerse achieves strong performance on sentiment and emotion tasks without requiring any specialized multitask learning module. Instead, it leverages the basic MLLM structure and the M2SE strategy to effectively analyze and reason about complex emotional contexts.
In addition, existing datasets often fail to provide the comprehensive and nuanced data necessary for training models that can effectively handle multiple tasks. This underscores the urgent need for a more robust dataset capable of supporting multitask learning in sentiment and emotion analysis. To address this gap, we introduce the Emotion Multitask (EMT) dataset, specifically designed to support MLLMs across five distinct tasks: multimodal sentiment analysis (MSA), emotion recognition (ER), facial expression recognition (FER), emotion reasoning inference (ERI), and emotion cause-pair extraction (ECPE) [33, 34]. By incorporating multiple tasks, the EMT dataset facilitates the development of models that can recognize, reason, and infer the causes of sentiment and emotion.
In summary, our main contributions are as follows:
  • We construct the EMT dataset. Each piece of data in this dataset contains queries for five tasks and the corresponding labels. The EMT dataset encourages the model to capture the relationship between sentiment and emotion from the perspective of different tasks, enabling the model to extract richer contextual information.
  • We introduce the M2SE strategy, which enables effective multitask learning by leveraging multistage training on diverse emotion and sentiment-related tasks. To the best of our knowledge, this is the first time the ECPE task has been introduced within the emotion domain as part of a multitask learning framework.
  • We develop the EmoVerse model. EmoVerse unifies tasks in the sentiment and emotion domains by employing the M2SE strategy. To ensure better adaptability to edge devices, we train the model with two different parameter configurations: 4B and 8B.
  • EmoVerse demonstrates outstanding performance across a wide range of tasks. On the CMU-MOSEI dataset for MSA[35], it achieves an Acc2 score of 88.51%. Similarly, on the MELD dataset for ER[36], it attains a weighted F1 score of 66.74. Additionally, ablation studies further validate the effectiveness of the M2SE strategy.
Figure 2: An example from the EMT dataset, which includes five tasks: sentiment analysis, emotion recognition, emotion reason inference, facial expression recognition, and emotion cause-pair extraction———along with its data construction process.
II Related Work
II-A Multimodal Large Language Models (MLLMs)
Multimodal Large Language Models (MLLMs) have gained significant attention due to their powerful reasoning capabilities. These models are typically trained for general-purpose tasks using a standard framework that includes modal encoders, a large language model (LLM), and a mapping layer (e.g., Linear [30], MLP [32, 37], or Q-Former [38]). While they excel in general tasks, MLLMs struggle with MSA and ER due to insufficient training on sentiment-specific datasets and emotion-related knowledge. Recent efforts have focused on improving MLLMs by incorporating multimodal emotion datasets and emotion reasoning tasks [39, 40]. However, these models still face challenges in leveraging contextual information effectively, particularly in sentiment analysis, where understanding the emotional nuances is crucial.
II-B Multitask Learning (MTL)
Multitask Learning (MTL) is a machine learning paradigm that enables models to simultaneously learn multiple related tasks, thereby allowing knowledge from one task to be transferred to others [41]. In affective computing, where multimodal inputs (e.g., text, audio, and visual frames) carry diverse and distinct information, MTL frameworks aim to leverage the synergies between tasks for improved performance. For instance, the MTMM-ES framework [42] jointly performs sentiment and emotion analysis, showing improved results over single-task models. UniMSE [11] unifies MSA and ER within a single framework. However, it relies solely on labeled data to establish the relationship between sentiment and emotion, without fully utilizing multimodal connections. M2Seq2Seq [43] proposes a multimodal multitask learning model based on the encoder-decoder architecture. Recently, Emotion-LLaMA [39] has shown promising results in ERC by incorporating multitask learning, including emotion cause inference. However, despite its success in ERC, Emotion-LLaMA still struggles with MSA tasks and fails to effectively synthesize contextual information for emotion inference.
Figure 3: Overview of the EmoVerse training process and model architecture. For different training stages, we assign tasks to the data based on their respective sampling rates, which are indicated in parentheses. The EmoVerse model is then trained using the assigned data. The model architecture include visual encoder, linear projector, and LLM.
III Methodology
This section presents our work, which consists of three key components: the construction of the EMT dataset, the formulation of the M2SE strategy, and the development of the EmoVerse model.
III-A EMT datasets
The EMT dataset is constructed by integrating data from the CMU-MOSEI, MELD, and ECF2.0[33] datasets to support five distinct tasks: MSA, ER, FER, ERI, and ECPC. Figure 2 illustrates a sample from the EMT dataset.
For the MSA and ER tasks, we use the original sentiment labels and emotion annotations from the source datasets as task labels. Due to the complexity of annotating the ECPE task, we refer to the annotation method of the ECF dataset and manually annotate a portion of the data. This is then combined with the existing ECF2.0 data to form the ECPE portion of the EMT dataset for further emotional causality analysis.
In the FER task, we adapt methods from the MERR dataset[39] by extracting peak frames based on the Action Units (AUs) calculated for each frame using the OpenFace tool. The composite score for each frame is determined by summing the AU values, as shown in the following equation:
$S_{f}=\sum_{i=1}^{n}\text{AU}_{i}(f)$ (1)
where $S_{f}$ is the composite score of frame $f$, and $\text{AU}_{i}(f)$ represents the Action Unit value for the $i$-th facial action unit at frame $f$. The peak frame with the highest score is selected to represent the most expressive moment. For videos featuring multiple characters, such as those in the MELD dataset, we improve the method by extracting peak frames separately for each character and then selecting the final peak frame by comparing the highest composite scores from all characters:
$S_{f_{\text{final}}}=\max\left(\{S_{f}^{(1)},S_{f}^{(2)},\dots,S_{f}^{(k)}\}\right)$ (2)
where $S_{f}^{(i)}$ is the composite score for character $i$ and $k$ represents the total number of characters in the frame. The final peak frame is the one with the highest composite score among all characters.
After selecting the peak frames, we extract the facial expression captions by combining the AUs from the peak frame with the predefined AU list corresponding to the emotion. Let $\text{AU}_{\text{frame}}(f)$ represent the set of AUs from the peak frame for a given emotion, and $\text{AU}_{\text{emotion}}(e)$ represent the predefined AU list for a specific emotion $e$. The common AUs are obtained by taking the intersection of the two sets:
$\text{AU}_{\text{common}}(f,e)=\text{AU}_{\text{frame}}(f)\cap\text{AU}_{\text{emotion}}(e)$ (3)
where $\text{AU}_{\text{common}}(f,e)$ is the set of AUs common to both the peak frame and the emotion-specific AU list. These common AUs are then mapped to textual information to generate the FER annotations for the task. This approach enables the model to link the facial expressions captured in the peak frames with the corresponding emotional state.
In the ERI task, we input the video into the Internvl2-8B  [^1] https://huggingface.co/OpenGVLab/InternVL2-8B model to generate detailed descriptions of both the characters and the scene. These descriptions allow the model to analyze emotional causes by considering facial expressions, speech, and environmental context. Finally, we use the LLaMA 3.1-8B  [^2] https://github.com/meta-llama/llama3 model to infer the underlying causes of the characters’ emotions by synthesizing the character and scene descriptions.
We perform a thorough preprocessing and quality control process on the raw data from the three datasets, systematically eliminating problematic or noisy data. Simultaneously, we conduct a manual curation of high-quality samples, following the aforementioned methodology. This involves refining the dataset to ensure accuracy and consistency, while extracting the necessary labels for FER and ERI. The final dataset, after this meticulous refinement, is outlined in Table I, including the number of tasks per category. Note that each data sample can be associated with multiple tasks. Table I below breaks down the number of tasks that can be used.
Additionally, we transform the dataset into a structured format, which includes three key elements: query, response, and images/videos. This format facilitates the instruction tuning of MLLMs.
TABLE I: The Number of Entries in EMT Dataset
Tasks Number of Entries Sentiment Analysis 25,859 Emotion Recognition 25,859 Facial Expression Recognition 15,870 Emotion Reason Inference 4,839 Emotion Cause-Pair Extraction 7,081 Total (Summed) 32,940
III-B EmoVerse
To validate the effectiveness of the M2SE strategy, we select the most widely used MLLM framework without introducing any multitask modifications to its architecture. This ensures that any observed performance improvements can be directly attributed to the proposed training strategy, rather than architectural changes. The proposed MLLM architecture, EmoVerse, as shown in Figure 3, consists of a visual encoder, a linear projector, and the LLM.
We train EmoVerse following M2SE strategy using the EMT datasets. For a data sample $D$, the input consists of the pair $\{v,t\}$, where $v$ represents the video and $t$ represents the text. First, the video $v$ is processed to extract visual tokens using a Vision Transformer (ViT)[44], producing a set of visual tokens $T_{v}$. These visual tokens $T_{v}$ are then passed through a linear transformation $\mathbf{W}_{v}$ to align the visual features with the feature space of the LLM. Mathematically, this can be expressed as:
$T^{\prime}_{v}=\mathbf{W}_{v}T_{v}$ (4)
where $T^{\prime}_{v}$ represents the aligned visual features.
Next, the text $t$, task identifiers, and instructions are tokenized and passed through the embedding layer of the LLM, producing the tokenized representation $T_{t}$. The tokenized visual features $T^{\prime}_{v}$ and the tokenized text $T_{t}$ are then concatenated to form the final input to the model:
$\mathbf{X}=[T^{\prime}_{v};T_{t}]$ (5)
where $\mathbf{X}$ is the combined input. The model processes this concatenated input and outputs the response $\hat{y}$, which corresponds to the task at hand:
$\hat{y}=\mathcal{M}(\mathbf{X})$ (6)
where $\mathcal{M}$ is the model and $\hat{y}$ is the model’s prediction or answer.
III-C Multi-Stage Multitask Sentiment and Emotion Instruction Tuning
III-C1 Stage 1: Multitask Pretraining
In the first phase of training, the model focuses on three key tasks: MSA, ER, and FER. We randomly select 15,000 samples from the EMT dataset for this phase, giving higher sampling rates to the MSA and FER tasks, while assigning a lower sampling rate to the ER task. The primary objective of this stage is to enable the model to recognize facial expressions and, through these expressions, capture the relationship between sentiment and emotion. Specifically, the FER task helps the model learn fine-grained facial features, which in turn improves performance in both the MSA and ER tasks by reinforcing the connection between emotional cues and sentiment.
III-C2 Stage 2: Multitask Reason Fine-Tuning
Following the first phase, where the model learns the fundamental relationship between sentiment and emotion as well as facial expression recognition, the second phase introduces four primary tasks: MSA, ER, ERI, and ECPE. For this phase, we select the remaining samples from the EMT dataset, randomly distributing them across the tasks. Notably, the MSA task is assigned a very small sampling rate, while the other three tasks—ER, ERI, and ECPE—are given higher and equal sampling rates. The ERI task enhances the model’s ability to reason about emotional causes by analyzing speech, facial expressions, and contextual information, building on the facial expression recognition learned in Stage 1. On the other hand, the ECPE task provides deeper insights into the causes of emotions by linking emotional responses with their underlying causes in conversations. Through the extraction of emotion cause-pairs, ECPE fosters better integration of multimodal and contextual information, helping the model understand the complex relationships between sentiment, emotion, and context. The details of Stage 1 and Stage 2 can be found in the Table II.
TABLE II: Stage 1 and Stage 2 with Tasks and Task Identifiers
Different Tasks Stages Task Identifier Stage 1 Stage 2 Sentiment Analysis <sentiment> Emotion Recognition <emotion> Facial Expression Recognition × <caption> Emotion Reason Inference × <reason> Emotion Cause-Pair Extraction × <emotion cause-pair>
IV Experiments
IV-A Experiments Setup
EmoVerse is trained solely on the EMT dataset without the addition of any external data, and it is tested on the following task-specific test sets.
IV-A1 Sentiment Analysis.
We use the CMU-MOSEI datasets for MSA. The CMU-MOSEI dataset is an extended version of the CMU-MOSI dataset. It encompasses 3228 videos from 1000 speakers, comprising a total of 23453 sentences. Similar to the CMU-MOSI dataset, it covers multiple emotional dimensions, employing the same annotation labeling approach. The results for Acc-2 calculated based on negative/non-negative (N/N) scheme, while the results on the right are calculated according to negative/positive (N/P) definition.
IV-A2 Emotion Recognition.
We use the MELD datasets for ER. The MELD dataset is an emotional dialogue dataset comprising 1433 dialogue segments and 13708 utterances from movies and TV shows. Each statement in the dialogue is labeled with one of seven emotions: anger, disgust, sadness, joy, neutral, surprise, and fear. Furthermore, the MELD data set provides annotations for emotional polarity (positive, negative, and neutral) for each utterance. We use Acc and weighted f1-sorce for evaluation.
IV-A3 Emotion Cause-pair Extraction.
We use the ECF2.0 datasets for ECPE. The ECF2.0 dataset is a comprehensive resource for emotion cause-pair extraction, comprising 1,715 conversations and 16,720 utterances. Each utterance in the dataset is annotated with emotion cause-pair information, which helps in understanding the emotional context and its causes. The dataset is divided into training and evaluation sets, with 1,374 conversations and 13,619 utterances in the training set, and 341 conversations and 3,101 utterances in the test set. We refer to previous evaluations and assess the ECF2.0 test set using the F1 Score and Weighted F1 Score.
IV-B Implementation Details
For the visual encoder, we utilize pre-train ViT with input images resized to $448\times 448$ pixels. During the fine-tuning process, the visual backbone is frozen, with the focus placed on training the linear projection layer. For LLM, we employ Phi-3-Mini[45] for EmoVerse-4B and InternLM-2.5-7B-Chat [^3] https://github.com/InternLM/InternLMfor EmoVerse-8B, both augmented with LoRA[46] for parameter-efficient fine-tuning.
In each stage of training, we train for 2 epochs with LoRA hyperparameters set to $r=8$ and $a=32$. The learning rate is set to $1\times 10^{-5}$, and we use a warm-up cosine schedule for the learning rate decay. For EmoVerse-4B, the training is conducted on 2 NVIDIA RTX 3090 GPUs, requiring approximately 48 hours. For EmoVerse-8B, the training is executed on a NVIDIA A100 GPU, taking around 55 hours.
IV-C Results
IV-C1 Performance comparison.
For MSA task, we compare our EmoVerse model with several SOTA baselines, including TFN[7], MulT[12], PMR[21], MISA[17], MMIM[14], Self-MM [13], BBFM[15], FDMER[16],CONKI[18], DMD[20], ConFEDE[19], and MRC-D3AE[27]. In ER task, we evaluate our model against baseline models with DialogueGCN [22], DialogueCRN[23], DAG-ERC[8], MM-DFN[24], EmoCaps[9], ERMC-DisGCN[25], MMGCN[26], GA2MIF[10], UniMSE[11], and FacialMMT[47].
Additionally, for both MSA and ER, we select advanced models capable of processing both images and videos. These include LLaVA-Video [48] and InternVL2 [49], both of which are designed to handle multimodal input efficiently. The performance comparison results, presented in Table III and Table IV, clearly demonstrate that EmoVerse achieves SOTA performance across both tasks, surpassing all the aforementioned baseline models in terms of accuracy and robustness.
For the ECPE task, we conduct tests solely on MLLMs, as they are better equipped to handle the complexity and multimodal nature of the task. The Table V shows the results of ECPE.
TABLE III: The performance comparison on CMU-MOSEI dataset.
Model Acc2 TFN - / 82.5 MulT - / 82.5 Self-MM 82.81 / 85.17 MMIM 82.24 / 85.97 BBFN - / 86.20 FDMER - / 86.10 MISA 83.60 / 85.50 ConKI 82.73 / 86.25 ConFEDE 81.65 / 85.82 DMD - / 86.60 PMR - / 83.30 MRC-D3AE 83.10 / 85.50 LLaVA-Video - / 82.82 InternVL2-4B - / 82.09 InternVL2-8B - / 85.30 EmoVerse-4B 85.51 / 87.80 EmoVerse-8B 85.93 / 88.51
TABLE IV: The performance comparison on MELD dataset.
Model Acc W.F1-Score DialogueGCN 59.46 58.10 DialogueCRN 60.73 58.39 DAG-ERC - 63.65 MM-DFN - 59.46 EmoCaps - 64.00 ERMC-DisGCN - 64.22 MMGCN 62.49 58.65 GA2MIF - 58.94 UniMSE 65.09 65.51 FacialMMT - 66.58 LLaVA-Video 23.87 22.65 InternVL2-4B 39.46 40.59 InternVL2-8B 42.18 44.13 EmoVerse-4B 67.28 66.58 EmoVerse-8B 67.78 66.74
TABLE V: The performance comparison on ECF2 dataset.
Model W.F1-Score F1-Score LLaVA-Video 22.59 22.78 InternVL2-4B 33.24 33.35 InternVL2-8B 40.57 40.51 EmoVerse-4B 72.30 72.21 EmoVerse-8B 73.54 73.62
IV-C2 Analysis of Emotion Reasoning.
To demonstrate the qualitative performance of EmoVerse, we present a detailed comparison of emotion reasoning results across various models. Figure 4 showcases the emotion reasoning outputs of five models: GPT-4o, Video-LLaMA2[50], LLaMA-Video, InternVL2, and EmoVerse. The video depicts a woman displaying a strong reaction of surprise, highlighting the models’ ability to infer emotional states from multimodal inputs.
GPT-4o and Video-LLaMA2 both classify the emotion in the video as ”confident” or ”assertive.” However, LLaMA-Video and InternVL2 correctly identify the emotion as ”surprise.” Despite this, these two models fail to fully integrate multimodal information for ERI. For instance, LLaMA-Video makes its inference solely based on facial expressions, while InternVL2 briefly acknowledges the facial expressions without providing a detailed explanation of the emotional cause, instead relying solely on textual cues for emotion reasoning. In contrast, EmoVerse effectively combines both multimodal video and text information to infer the cause of the emotion, yielding a correct and comprehensive emotional interpretation.
Figure 4: Comparison of different models on ERI. It can be seen that both GPT-4o and Video-LLaMA2 make errors in ER. While Llava-Video and InternVL2 correctly classify the emotions, their descriptions of the reasons behind the emotions are incomplete. In contrast, EmoVerse not only correctly classifies the emotions but also provides a comprehensive description of the reasons behind them.
IV-D Ablation Evaluation
We use EmoVerse-4B to do the following ablation experiments.
IV-D1 Comparison of the M2SE Strategy with Single-Task Fine-Tuning.
To evaluate the impact of the M2SE strategy within the same model framework and compare its performance across various tasks, we conduct an ablation study. As shown in Table VI, the tasks in parentheses represent the fine-tuning tasks, with each task corresponding to the same datasets used in our evaluation. ”Z.S.” stands for zero-shot. The M2SE strategy is trained using the EMT dataset. The results demonstrate that, although the performance on some metrics may not exceed that of single-task learning, EmoVerse trained with the M2SE strategy consistently performs better overall compared to other models.
TABLE VI: Comparison of the M2SE Strategy with Single-Task Fine-Tuning
Model CMU-MOSEI MELD ECF2.0 Acc2 Acc W.F1 F1 W.F1 EmoVerse(Z.S.) 82.09 / 83.87 39.46 40.59 33.24 33.35 EmoVerse(MSA) 83.96 / 88.16 48.12 31.27 41.67 41.14 EmoVerse(ER) 78.99 / 80.17 66.21 65.71 40.61 38.23 EmoVerse(EPCE) 83.27 / 85.85 32.38 33.43 72.69 72.59 EmoVerse(M2SE) 85.51 / 87.80 67.78 66.74 72.30 72.21
IV-D2 Effect of the FER Task.
To explore the contribution of the FER task to MSA, we conduct an ablation experiment to compare the effects of increasing and decreasing the FER sampling ratio on MSA performance, using the same training data volume in the first stage. For a low sampling ratio, we denote it as ”lf,” and for a high sampling ratio, we denote it as ”hf.” As shown in Table VII, under the same data volume, increasing the FER task ratio—despite the model performing fewer MSA tasks—still leads to stronger learning performance. This indicates that even with fewer MSA samples, the contribution of FER significantly enhances the model’s ability to perform sentiment analysis.
TABLE VII: Effect of the FER Task.
Model Acc2 EmoVerse (lf) 77.09 / 87.47 EmoVerse (hf) 82.61 / 87.50
IV-D3 Effect of the Choice of Tasks in the Second Stage.
To investigate how task configurations impact model performance and better understand cross-task interactions, we conduct an ablation study in the second stage with three setups: ”t1” (only ER task), ”t3” (ER, ERI, and ECPE tasks), and ”t4” (all four tasks). Results, as shown in Table VIII, show that focusing solely on ER (t1) improves ER performance but weakens MSA and emotion reasoning tasks. Adding ERI and ECPE (t3) slightly lowers MSA and ER performance but enhances the model’s emotion reasoning ability. When MSA is reintroduced with a small sampling rate (t4), the model effectively integrates cross-task knowledge, improving performance across all tasks and surpassing first-stage MSA results. This demonstrates the importance of balanced task integration for holistic model improvement.
TABLE VIII: Effect of the Distribution of Tasks on EmoVerse.
Model CMU-MOSEI MELD ECF2.0
Acc2 Acc W.F1 F1 W.F1
EmoVerse (f) 82.61 / 87.50 - - - -
EmoVerse (t1) 81.68 / 87.20 66.74 66.24 25.40 26.69
EmoVerse (t3) 79.0 / 86.78 66.97 65.67 69.85 69.95
EmoVerse (t4) 85.51 / 87.80 67.78 66.74 72.30 72.21
V Conclusion
In this paper, we propose the M2SE strategy to guide the training process, enabling the model to learn from multiple sentiment- and emotion-related tasks across different stages. To support this, we construct the EMT dataset, which includes a large number of annotated samples spanning a wide range of tasks. Additionally, we introduce EmoVerse, a unified MLLM designed to enhance emotion perception by effectively integrating both visual and textual data. By leveraging multitask and multistage learning, EmoVerse unifies sentiment and emotion tasks, achieving superior performance in MSA, ER, and EPCE. It outperforms existing models on benchmark datasets such as CMU-MOSEI and MELD. The experimental results highlight the effectiveness of the M2SE strategy in enabling the model to capture and reason about complex emotional contexts, showcasing its potential to drive advancements in emotion-aware AI applications.
Acknowledgments
This work was supported by the Key R&D Program of Shandong Province, China (Major Scientific and Technological Innovation Project) (NO.2022CXGC010504) and National Natural Science Foundation of China under Grant 61301253.